OpenStudy (anonymous):

Finding the slope of a line: A(2,6) C(6,-2)

OpenStudy (anonymous):

Slope = change in y/change in x. Therefore change in y=-2-6=-8 and change in x=6-2=4. So we get slope=-8/4=-2

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

sry -2